# Quasi-alternating surgeries

### v1

@@ 2.2. Symmetry groups of hyperbolic knots
Let $M$ be a $3$-manifold. The symmetry group of $M$ is defined to be the mapping class group of $M$, i.e. the group of orientation-preserving diffeomorphisms of $M$ that fix the boundary pointwise. If $M$ is hyperbolic then the symmetry group of $M$ is isomorphic to its isometry group [Gab97, GMT03]. The symmetry group of a hyperbolic knot (i.e. the symmetry group of its complement) is known to be a subgroup of a dihedral group [HTW98] and thus it is $\mathbb{Z}_{n}$, $D_{n}$, or $\mathbb{Z}_{2}^{2}$.
An orientation-preserving involution on $S^{3}$ is either the antipodal map which has no fixed set or a $\pi$-rotation about the fixed set of a great circle [Wal69]. A free involution on a knot $K$ is an involution of $(S^{3},K)$ that acts without fixed points on $K$. These come in two types depending on whether or not the action has fixed points in $S^{3}$. A strong inversion on a knot $K$ is an involution of $(S^{3},K)$ with a circle fixed set that meets $K$ in two points. In both situations, $K$ has a tubular neighborhood that is equivariant. Observe that the quotient of a strong involution of $K$, restricted to the exterior of an equivariant tubular neighborhood of $K$, is a tangle exterior for $K$ [Mon75].
@@
Let $f$ be an orientation-preserving involution on the exterior $M_{K}=S^{3}\backslash\mathcal{N}(K)$ of $K$. Then $f$ must preserve the slope of the longitude of $K$ in $\partial\mathcal{N}(K)$: $f(\lambda)=\pm\lambda$. Furthermore, due to the solution of the knot complement problem [GL89], $f$ must also preserve the slope of the meridian of $M$: $f(\mu)=\pm\mu$. Since $f$ is orientation-preserving on $M_{K}$, the restriction of $f$ to $\partial\mathcal{N}(K)$ is either the identity $I$ or its negation $-I$. In both cases, this map $f$ extends to an involution of $S^{3}$ that preserves $K$. The former gives a free involution while the latter gives a strong inversion. ∎
@@ 2.3. Double branched covers
Let $M$ be a closed $3$-manifold with a smooth involution $\varphi\colon M\rightarrow M$. We can take the quotient under $\varphi$, i.e. we identify points $p$ and $\varphi(p)$. Then the quotient map $M\rightarrow M/\varphi$ will be a (branched or unbranched) $2$-fold covering map with (possibly empty) branching set $B_{\varphi}$. For this, we also write $M=\operatorname{DBC}(B_{\varphi})$. It will be an unbranched covering if and only if the involution acts freely. The following lemma says that the quotient manifold and its branching set only depend on the mapping class group of $\varphi$.
@@
Suppose $M$ is a closed hyperbolic $3$-manifold with two smooth involutions $\varphi_{1},\varphi_{2}\colon M\rightarrow M$ that are isotopic as diffeomorphisms. Then $(M/\varphi_{1},B_{\varphi_{1}})$ is diffeomorphic to $(M/\varphi_{2},B_{\varphi_{2}})$.
@@
We see $M/\varphi_{i}$ as an orbifold $\mathcal{O}_{i}$. Then the quotient map $M\rightarrow\mathcal{O}_{i}$ is a $2$-fold orbifold covering map. Since $M$ is hyperbolic, the orbifold universal covering of $\mathcal{O}_{i}$ is hyperbolic, and thus $\mathcal{O}_{i}$ is irreducible and atoroidal. Then the orbifold geometrization theorem [BLP05, Corollary 1.2] implies that $\mathcal{O}_{i}$ is hyperbolic and the deck transformation $\varphi_{i}$ is an isometry for some hyperbolic metric $g_{i}$ on $M$. Now Mostow rigidity implies that there exists an isometry $f\colon M\rightarrow M$ such that $f^{-1}\varphi_{2}f=\varphi_{1}$. But then $\mathcal{O}_{1}\ni[p]\mapsto[\varphi_{2}f(p)]\in\mathcal{O}_{2}$ is a diffeomorphism respecting the branching sets. ∎
In this work, the only non-trivial symmetry groups of closed hyperbolic $3$-manifolds that we need to consider are the symmetry groups $\mathbb{Z}_{2}$ and $\mathbb{Z}_{2}^{2}$. In the former case, there is a unique involution. In the latter case, the symmetry group is generated by two different involutions $f$ and $g$. The involution $f\circ g$ will always be different from $f$ and $g$ since it represents a different element in the symmetry group. (However, $g\circ f$ is isotopic to $f\circ g$.) If we present our closed hyperbolic $3$-manifold via a surgery diagram, we can often explicitly determine each involution via symmetries of the surgery link and then explicitly determine the corresponding quotient and branch locus there. For a more detailed discussion see [Auc14].
@@ 2.4. Length of geodesics after surgery and symmetry-exceptional slopes
In this section, building upon [FPS22], we will give an explicit bound $C$ such that any surgery on a hyperbolic knot along a slope whose length is larger than $C$ yields again a hyperbolic manifold with the same symmetry group as the knot. We first set up the notation.
Let $M$ be a complete hyperbolic $3$-manifold with finite volume and one cusp. We denote by $\operatorname{systole}(M)$ the length of the shortest geodesic of $M$. The (normalized) length $L(r)$ of a slope $r$ on $\partial M$ is
$L(r)=\frac{\operatorname{length}(r)}{\sqrt{\operatorname{area}(\partial M)}}.$
We denote by $M(r)$ the Dehn filling of $M$ along the slope $r$. If the slope $r$ is in the hyperbolic Dehn surgery space (so that $M(r)$ has a hyperbolic structure obtained from deforming the one on $M$) then we write $\gamma$ for the geodesic core of the newly glued-in solid torus in $M(r)$. From Thurston’s hyperbolic Dehn surgery theorem [Thu79], it follows that $\gamma$ will be the shortest geodesic in $M(r)$ provided that $L(r)$ is sufficiently large.
The following theorem, which is a slight reformulation of Theorem 7.28 in [FPS22], hands us an explicit such bound on the length of the slope.
Theorem 2.4 (Futer–Purcell–Schleimer [FPS22])
@@
Note that $L(r)\geq C(M)>6$, and thus the manifold $M(r)$ is necessarily hyperbolic by [Ago00, Lac00], cf. [BH96]. Note that the $6$-theorem is in terms of non-normalized length, so we are implicitly using the fact that there always exists a cusp of area at least one, see Theorem 1.2 in [GHM+21].
@@
When $L(r)\geq C(M)$, Theorem 2.4 says that (in addition to $M(r)$ being the hyperbolic manifold obtained by deforming the hyperbolic structure on $M$) the core $\gamma$ of the newly glued-in solid torus is the shortest geodesic in $M(r)$. It follows that any non-trivial element in the symmetry group of $M(r)$ has to send $\gamma$ to itself and thus restricts to a symmetry of the knot complement $M$. ∎
@@ 3. Algorithms and Strategies
In this section, we discuss certain strategies and algorithms that we use later to prove our main theorems. We will apply these strategies to certain hyperbolic L-space knots whose symmetry groups are either trivial or generated by a single strong inversion. It is conjectured that this covers all symmetries of hyperbolic L-space knots [BR23, Conjecture 1.7]. Thus we restrict the strategies to such knots whenever that simplifies the discussion.
@@
    Using the cusp shape of $K$, make a list of short slopes of length at most $6$. Every exceptional slope is contained in this finite list by the 6-Theorem [Ago00, Lac00].
@@
      If no hyperbolic structure is found, check with Regina and Dunfield’s recognition code [Dun20b] for classification of non-hyperbolic type. If this recognizes $K(r)$ as non-hyperbolic, $r$ is an exceptional slope.
@@
Dunfield employs such a strategy to classify exceptional slopes of census knots [Dun20b]. Since it is algorithmically decidable if a manifold is hyperbolic [Man02], this strategy yields an algorithm that is always guaranteed to terminate.
@@
    Using the cusp shape and systole computations of $K$, make a list of short slopes of length less than the FPS bound, Theorem 2.4. By Corollary 2.5 this finite list contains all symmetry-exceptional slopes.
@@
Strategy 3.5 (Branch locus search)
@@
    Iterate through links in 3-manifolds by taking links in the HTW table and performing Dehn fillings on the components, where some subset of the components receive the trivial $1/0$ filling. This gives a pair $(M,B)$ where $M$ is the surgered manifold and $B$ is the link in $M$ formed by the components with $1/0$ filling. Take all double covers of $M$ branched along $B$ [^2] Note that there might be more than one or no double branched cover depending on the algebraic topology of the link complement. For each cover $C$:
@@
As explained in Remark 3.4, Step $(1)$ of this algorithm is theoretically not always guaranteed to terminate (since it rests on symmetry group computations) but works well in practice and yields verified results. On the other hand, the rest of the algorithm is theoretically verified to terminate since the decision problem for $3$-manifolds is solved [Kup19]. But in practice, some of the branching sets may be too complicated to find in a reasonable time. Nevertheless, in most of our examples the branching sets were sufficiently simple that Strategy 3.5 terminated relatively quickly.
@@
    Using the Branch Locus Search 3.5, identify for some integer $n\in\mathbb{Z}$ a triple of links $B_{n}$, $B_{n+1}$ and $B_{n+1/2}$ in $S^{3}$ such that
@@
      If so, check if changing the crossing at $c$ unknots $B_{n+1/2}$ using the unknot recognition algorithm [BBPea].
@@
        Let $\delta$ be a crossing disk for the crossing $c$. Form the double branched cover $N$ of $S^{3}-\partial\delta$ along $B$.
@@
        The Conway sphere $\partial N(\delta)$ lifts to a torus $F$ in $N$ that bounds a knot manifold to one side and a pair-of-pants ×$\times S^{1}$ to the other side. SnapPy [CDGW] can search for the splitting torus $F$ in $N$ as a normal surface and perform operations on it.
@@
    The brute force approach of Strategy 3.7 is not guaranteed to succeed and can certainly be improved in speed and general applicability. Nevertheless, Strategy 3.7 works efficiently and successfully to identify a tangle exterior for each strongly invertible census knot. Moreover, because the Branch Locus Search (Strategy 3.5) identifies branch loci with low crossing number, the tangle exteriors that Strategy 3.7 identify also have low crossing number and their integral rational tangle fillings remain rather simple.
@@
    There is an algorithm to identify a tangle exterior for a strongly invertible knot, though it is impractical. Fix an enumeration by crossing number of all reduced diagrams of 2-strand tangle exteriors, and then inductively check if their double branched covers are diffeomorphic to the exterior of the knot at hand. This last step uses the solution to the decision problem for diffeomorphisms of 3-manifolds with torus boundary [Kup19].
@@
    If $K$ is strongly invertible, then every slope is a dbc slope by the Montesinos trick [Mon75]. For getting all branching sets we run the tangle exterior search 3.7 to find the tangle exterior $T_{K}$ of $K$. Then a branching set of $K(r)$ is given by the rational tangle filling $T_{K}(r)$.
@@
      If $r$ is exceptional, then use either a Seifert fibred structure on $K(r)$ or the Equivariant Torus Theorem [Hol91] for ad hoc determination.
@@
      If $r$ is symmetry-exceptional, perform the Branch Locus Search 3.5 to account for every involution in $\mathrm{Sym}(K(r))$. (Eg. if the involutions generate a $\mathbb{Z}_{2}^{2}$ subgroup, find three distinct branch loci.)
@@
Strategy 3.9 is not verified to terminate since it rests on symmetry group computations.
@@
Begin with lists $\mathcal{QA}$ of known quasi-alternating and $\mathcal{NQA}$ of known non-quasi-alternating links. Initially, we use Jablan’s classification of prime quasi-alternating knots with at most 12 crossings [Jab14] to construct these two lists. Each time this strategy is run, these lists grow.
@@
Since there exist links that are thin but not quasi-alternating [Gre10], Strategy 3.11 will in general not terminate. In fact, it remains unknown if it is algorithmically decidable if a link is quasi-alternating or not, see Question 9.6. The same holds true for Strategy 3.13 below.
@@
      Use Strategy 3.5 to search for a branch locus $L$ of $K(r)$.

### v2

@@ 2.2.2. Formal L-spaces
Remark 2.2
Formal L-spaces can be seen as the $3$-manifold analog of quasi-alternating links. In both cases, one obtains the corresponding class by starting from the simplest object (the unknot in the link setting and $S^{3}$ in the 3-manifold setting) and then closing under a recursive operation governed by additivity of determinant or homology in a triad. For quasi-alternating links, the skein triple $(L,L_{0},L_{1})$ at a crossing plays the role of a triad with the determinant relation $\det(L)=\det(L_{0})+\det(L_{1})$ mirroring the homological relation
$|H_{1}(Y;\mathbb{Z})|=|H_{1}(Y_{0};\mathbb{Z})|+|H_{1}(Y_{1};\mathbb{Z})|$
in a Dehn filling triad. Since the determinant of a link gives the order of the first homology of its double branched cover it follows that the double branched cover of a quasi- alternating link is a formal L-space. Thus, the definition of formal L-spaces can be viewed as a natural 3-manifold analogue of quasi-alternating links.
@@ 2.3. Symmetry groups of hyperbolic knots
Let $M$ be a $3$-manifold. The symmetry group of $M$ is defined to be the mapping class group of $M$, i.e. the group of orientation-preserving diffeomorphisms of $M$ that fix the boundary pointwise up to isotopy that fixes the boundary pointwise. If $M$ is hyperbolic then the symmetry group of $M$ is isomorphic to its isometry group [Gabai_DiffMCG, Gabai_DiffMCG2]. The symmetry group of a hyperbolic knot (i.e. the symmetry group of its complement) is known to be a subgroup of a dihedral group [HTW_table] and thus it is $\mathbb{Z}_{n}$, $D_{n}$, or $\mathbb{Z}_{2}^{2}$.
An orientation-preserving involution on $S^{3}$ is either the antipodal map, which has no fixed set, or a $\pi$-rotation about the fixed set of a great circle [Waldhausen_involutions]. A free involution on a knot $K$ is an involution of $(S^{3},K)$ that acts without fixed points on $K$. These come in two types depending on whether or not the action has fixed points in $S^{3}$. A strong inversion on a knot $K$ is an involution of $(S^{3},K)$ with a circle fixed set that meets $K$ in two points. In both situations, $K$ has a tubular neighborhood that is equivariant. Observe that the quotient of a strong involution of $K$, restricted to the exterior of an equivariant tubular neighborhood of $K$, is a tangle exterior for $K$ [Montesinos_trick].
@@
Let $f$ be an orientation-preserving involution on the exterior $M_{K}=S^{3}\backslash\mathcal{N}(K)$ of $K$. Then $f$ must preserve the slope of the longitude of $K$ in $\partial\mathcal{N}(K)$: $f(\lambda)=\pm\lambda$. Furthermore, due to the solution of the knot complement problem [GL89], $f$ must also preserve the slope of the meridian of $M$: $f(\mu)=\pm\mu$. Since $f$ is orientation-preserving on $M_{K}$, the restriction of $f$ to $\partial\mathcal{N}(K)$ is either the identity $I$ or its negation I-I. In both cases, this map $f$ extends to an involution of $S^{3}$ that preserves $K$. The former gives a free involution while the latter gives a strong inversion. ∎
@@ 2.4. Double branched covers
Let $M$ be a closed $3$-manifold with a smooth orientation-preserving involution $\varphi\colon M\rightarrow M$. We can take the quotient under $\varphi$, i.e. we identify points $p$ and $\varphi(p)$. Then the quotient map $M\rightarrow M/\varphi$ will be a (branched or unbranched) $2$-fold covering map with (possibly empty) branching set $B_{\varphi}$. For this, we also write $M=\operatorname{DBC}(B_{\varphi})$. It will be an unbranched covering if and only if the involution acts freely. The following lemma says that the quotient manifold and its branching set only depend on the isotopy class of $\varphi$ in the mapping class group of $M$.
@@
Suppose $M$ is a closed hyperbolic $3$-manifold with two smooth orientation-preserving involutions $\varphi_{1},\varphi_{2}\colon M\rightarrow M$ that are isotopic as diffeomorphisms. Then $(M/\varphi_{1},B_{\varphi_{1}})$ is diffeomorphic to $(M/\varphi_{2},B_{\varphi_{2}})$.
@@
We see $M/\varphi_{i}$ as an orbifold $\mathcal{O}_{i}$. Then the quotient map $M\rightarrow\mathcal{O}_{i}$ is a $2$-fold orbifold covering map. Since $M$ is hyperbolic, the orbifold universal covering of $\mathcal{O}_{i}$ is hyperbolic, and thus $\mathcal{O}_{i}$ is irreducible and atoroidal. Then the orbifold geometrization theorem [BLP_annals, Corollary 1.2] implies that $\mathcal{O}_{i}$ is hyperbolic and the deck transformation $\varphi_{i}$ is an isometry for some hyperbolic metric $g_{i}$ on $M$. Now Mostow rigidity implies that there exists an isometry $f\colon(M,g_{1})\rightarrow(M,g_{2})$ such that $f^{-1}\varphi_{2}f=\varphi_{1}$. But then $\mathcal{O}_{1}\ni[p]\mapsto[\varphi_{2}f(p)]\in\mathcal{O}_{2}$ is a diffeomorphism respecting the branching sets. ∎
In this work, the only non-trivial symmetry groups of closed hyperbolic $3$-manifolds that we need to consider are the symmetry groups $\mathbb{Z}_{2}$ and $\mathbb{Z}_{2}^{2}$. In the former case, there is a unique involution. In the latter case, the symmetry group is generated by two different involutions $f$ and $g$. The involution $f\circ g$ will always be different from $f$ and $g$ since it represents a different element in the symmetry group. (However, $g\circ f$ is isotopic to $f\circ g$.) If we present our closed hyperbolic $3$-manifold via a surgery diagram, we can often explicitly determine each involution via symmetries of the surgery link and then explicitly determine the corresponding quotient and branch set there. For a more detailed discussion see [Au14].
@@ 2.5. Length of geodesics after surgery and symmetry-exceptional slopes
In this section, building upon [FPS19], we will give an explicit bound $C$ such that any surgery on a hyperbolic knot along a slope whose length is larger than $C$ yields again a hyperbolic manifold with the same symmetry group as the knot. We first set up the notation.
Let $M$ be a complete hyperbolic $3$-manifold with finite volume and one cusp. We denote by $\operatorname{systole}(M)$ the length of the shortest geodesic of $M$. There is a (normalized) length $L(r)$ associated to each slope $r$ on $\partial M$. This is obtained by taking a horospherical neighbourhood $N$ of the cusp with the natural inherited Euclidean metric on $N$ and setting
$L(r)=\frac{\operatorname{length}(r)}{\sqrt{\operatorname{area}(\partial N)}}.$
where $\operatorname{length}(r)$ is the minimal length of all curves representing $r$ on $\partial N$. Since changing the horospherical neighbourhood $N$ has the effect of rescaling the Euclidean metric on $\partial N$, this $L(r)$ is independent of the choice of $N$.
We denote by $M(r)$ the Dehn filling of $M$ along the slope $r$. If the slope $r$ is in the hyperbolic Dehn surgery space (so that $M(r)$ has a hyperbolic structure obtained from deforming the one on $M$) then we write $\gamma$ for the geodesic core of the newly glued-in solid torus in $M(r)$. From Thurston’s hyperbolic Dehn surgery theorem [Th79], it follows that $\gamma$ will be the shortest geodesic in $M(r)$ provided that $L(r)$ is sufficiently large.
The following theorem, which is a slight reformulation of Theorem 7.28 in [FPS19], hands us an explicit such bound on the length of the slope.
Theorem 2.5 (Futer–Purcell–Schleimer [FPS19])
@@
Note that $L(r)\geq C(M)>6$, and thus the manifold $M(r)$ is necessarily hyperbolic by [Ag00, La00], cf. [BH96]. Note that the $6$-theorem is in terms of non-normalized length, so we are implicitly using the fact that there always exists a cusp of area at least one, see Theorem 1.2 in [gabai2021hyperbolic].
@@
When $L(r)\geq C(M)$, Theorem 2.5 says that (in addition to $M(r)$ being the hyperbolic manifold obtained by deforming the hyperbolic structure on $M$) the core $\gamma$ of the newly glued-in solid torus is the shortest geodesic in $M(r)$. It follows that any element in the symmetry group of $M(r)$ has to send $\gamma$ to itself and thus restricts to a symmetry of the knot complement $M$. ∎
@@ 3. Algorithms and Strategies
In this section, we discuss certain strategies and algorithms that we use later to prove our main theorems. We will apply these strategies to certain hyperbolic L-space knots whose symmetry groups are either trivial or generated by a single strong inversion. It is conjectured that this covers all symmetries of hyperbolic L-space knots [L_space_periodic_conj, Conjecture 1.7]. Thus we restrict the strategies to such knots whenever that simplifies the discussion.
@@
    Using the cusp shape of $K$, make a list of short slopes of length at most $6$. Every exceptional slope is contained in this finite list by the 6-Theorem [Ag00, La00].
@@
      If no hyperbolic structure is found, check with Regina and Dunfield’s recognition code [Du18] for classification of non-hyperbolic type. If this recognizes $K(r)$ as non-hyperbolic, $r$ is an exceptional slope.
@@
Dunfield employs such a strategy to classify exceptional slopes of census knots [Du18]. Since it is algorithmically decidable if a manifold is hyperbolic [decide_hyperbolic], this strategy yields an algorithm that is always guaranteed to terminate.
@@
    Using the cusp shape and systole computations of $K$, make a list of short slopes of length less than the FPS bound, defined in Theorem 2.5. By Corollary 2.6 this finite list contains all symmetry-exceptional slopes.
@@
Strategy 3.5 (Branch set search)
@@
    Iterate through links in 3-manifolds by taking links in the HTW table and performing Dehn fillings on the components, where some subset of the components receive the trivial $1/0$ filling, measured with respect to the Seifert longitude in $S^{3}$. This gives a pair $(M,B)$ where $M$ is the surgered manifold and $B$ is the link in $M$ formed by the components with $1/0$ filling. Take all double covers of $M$ branched along BB. [^6] Note that there might be more than one or no double branched cover depending on the algebraic topology of the link complement. For each cover $C$:
@@
As explained in Remark 3.4, Step $(1)$ of this algorithm is theoretically not always guaranteed to terminate (since it rests on symmetry group computations) but works well in practice and yields verified results. On the other hand, the rest of the algorithm is theoretically verified to terminate since the decision problem for $3$-manifolds is solved [Kuperberg_Decision_prob]. But in practice, some of the branching sets may be too complicated to find in a reasonable time. Nevertheless, in most of our examples, the branching sets were sufficiently simple that Strategy 3.5 terminated relatively quickly.
Note that Step $(3)$ and $(4)$ will in theory always terminate and return all branching sets, since every link $B$ in a $3$-manifold $M$ admits a surgery description along a link in $S^{3}$. However, in practice we iterate only through links in the HTW table with a small subset of chosen surgery coefficients.
@@
    Using the branch set search 3.5, identify for some integer $n\in\mathbb{Z}$ a triple of links $B_{n}$, $B_{n+1}$ and $B_{n+1/2}$ in $S^{3}$ such that
@@
      If so, check if changing the crossing at $c$ unknots $B_{n+1/2}$ using the unknot recognition algorithm [BBP+]. If this is the case, we expect that the exterior of $c$ is the tangle exterior of $K$, since four of its tangle fillings give the correct branching sets.
@@
        Let $\delta$ be a crossing disk for the crossing $c$, i.e. an embedded disk $\delta\subset S^{3}$ transversally intersecting the link $B$ at exactly two points in its interior that projects to an arc in the diagram plane meeting $D$ only at the crossing $c$. Form the double branched cover $N$ of $S^{3}-\partial\delta$ along $B$.
@@
        The Conway sphere $\partial N(\delta)$ lifts to a torus $F$ in $N$ that bounds a knot manifold to one side and a pair-of-pants $\times S^{1}$ to the other side. SnapPy [CDGW] can search for the splitting torus $F$ in $N$ as a normal surface and perform operations on it.
@@
    The brute force approach of Strategy 3.7 is not guaranteed to succeed and can certainly be improved in speed and general applicability. For example we could try different integers $n$ in Step $(1)$ or different diagrams $D$ in Step $(2)$. Nevertheless, Strategy 3.7 works efficiently and successfully to identify a tangle exterior for each strongly invertible census knot. Moreover, because the branch set search (Strategy 3.5) identifies branch loci with low crossing number, the tangle exteriors that Strategy 3.7 identify also have low crossing number and their integral rational tangle fillings remain rather simple.
@@
    There is an algorithm to identify a tangle exterior for a strongly invertible knot, though it is impractical. Fix an enumeration by crossing number of all reduced diagrams of 2-strand tangle exteriors, and then inductively check if their double branched covers are diffeomorphic to the exterior of the knot at hand. This last step uses the solution to the decision problem for diffeomorphisms of 3-manifolds with torus boundary [Kuperberg_Decision_prob].
@@
    If $K$ is strongly invertible, then every slope is a dbc slope by the Montesinos trick [Montesinos_trick]. For getting all branching sets we run the tangle exterior search 3.7 to find the tangle exterior $T_{K}$ of $K$. Then a branching set of $K(r)$ is given by the rational tangle filling $T_{K}(r)$.
@@
      If $r$ is exceptional, then use either a Seifert fibred structure on $K(r)$ or the Equivariant Torus Theorem [equivarianttorustheorem] for ad hoc determination. Further discussion on how this can be done in practice can be found in Section 4.1 in [BKM_thin].
@@
      If $r$ is symmetry-exceptional, perform the branch set search 3.5 to account for every involution in $\mathrm{Sym}(K(r))$. (Eg. if the involutions generate a $\mathbb{Z}_{2}^{2}$ subgroup, find three distinct branch loci.)
@@
Strategy 3.9 is not verified to terminate since it rests on symmetry group computations, see Remark 3.4.
@@
Begin with lists $\mathcal{QA}$ of known quasi-alternating and $\mathcal{NQA}$ of known non-quasi-alternating links. Initially, we use Jablan’s classification of prime quasi-alternating knots with at most 12 crossings [Ja14] to construct these two lists. Each time this strategy is run, these lists grow.
@@
Since there exist links that are thin but not quasi-alternating [Greene_thin_QA], Strategy 3.11 will in general not terminate. In fact, it remains unknown if it is algorithmically decidable if a link is quasi-alternating or not, see Question 9.6. The same holds true for Strategy 3.13 below.
@@
      Use Strategy 3.5 to search for a branch set $L$ of $K(r)$.